My Wonderful Article in IEEE Format

Pablo L. Sala Computer Science Department University of Toronto Toronto, ON, M5R 3L1

Abstract

This is the abstract of my paper. It must fit within the size allowed, which is about 3 inches, including section title, which is 11 point bold font. If you don’t want the text in italics, simply remove the ‘em’ command and the curly braces which bound the abstract text. If you have em commands within an italicized abstract, the text will come out as normal (non-italicized) text.

1. Introduction

Here is my introduction text. The section heading is in a large bold font and all sections and subsections are numbered.

Type your main text in 10-point Times, single-spaced at 12pt. Do not use double-spacing. All paragraphs should be indented 1 pica (approximately 1/6- or 0.17-inch or 0.422 cm). Be sure your text is fully justified—that is, flush left and flush right. Please do not place any additional blank lines between paragraphs.

1.1. Previous Work

There are various bibliographic and citation schemes available in LaTex, but we choose to use the simplest one in this example. Throughout I may cite references of the form [?] or [?], and LaTeX will keep track of numbering. The numbers are based on the order you place them in the bibliography, not the order they appear in the text. They should (I believe) be in alphabetical order. LaTex will put square brackets about the number within the text of your paper. For those of you new to LaTex, you may have to run the latex process twice to allow all references to be resolved. You will get a warning about a missing .aux file. Just rerun latex and it will be ok.

2. Summary and Conclusions

This template will get you through the minimum article, i.e., with no figures or equations. To include those, please refer to your LaTeX manual and the IEEE publications guidelines. However, for a vision conference you will probably want the following equation somewhere:

g(x) = \frac{1}{\sqrt{2π}σ} e^{−x^{2}/2σ^{2}}

Good Luck!

Acknowledgments

This is how to do an unnumbered subsection. For submission, there should be no acknowledgments as this could lead to identification of the author.

References

[1] I. M. Author, “Some Related Article I Wrote,” Some Fine Journal, Vol. 17, pp. 1-100, 1987.

[2] A. N. Expert, A Book He Wrote, His Publisher, 1989.